DO NOT ANSWER IF YOU ARE UNSKILLED IN THIS AREA!

Exercise 14-43 Algo

In a simple linear regression based on 58 observations, it is found that SSE = 2,622 and SST = 26,815.

a. Calculate s2ese2 and se. (Round your answers to 2 decimal places.)

s2e
se


b. Calculate the coefficient of determination R2. (Round your answer to 4 decimal places.)

Coefficient of Determination
Answer #1

Given that, n= 58

SSE = 2622 and SST = 26815

=> SSR = SST - SSE = 26815 - 2622 = 24193

a)

S_e^2= \frac { SSE }{n-2}= \frac { 2622}{58 -2}=\frac {2622 }{56} = 46.82

S_e = \sqrt { \frac { SSE }{n-2}}= \sqrt {\frac { 2622}{58 -2}} = \sqrt {\frac {2622 }{56} } = 6.84

b) Coefficient of determination is,

R2 = SSR / SST = 24193 / 26815 = 0.9022

=> R2 = 0.9022